OpenStudy (anonymous):

Subtract and simplify: x^2/x+4 - 16/x+4

OpenStudy (anonymous):

@Zale101 @charlotte123

OpenStudy (gorv):

(x^2-16)/(x+4)=(x-4)*(x+4)/(x-4)=x+4